# Likelihood and appearance of life beyond the Earth: An astronomical perspective

### v1

@@
Abstract
As of 2023, over 5500 planets are known to orbit stars other than our Sun. We can measure their sizes and orbital periods, infer their masses and temperatures, and constrain their compositions. Based on these data, about 1% of extrasolar planets are potentially habitable for life as we know it, implying that of the billions of planets in our Galaxy, some may actually be inhabited, at least by microbes. However, recognizing signs of alien life forms is a major challenge for current technology, because of the wide range of conditions on extrasolar planets, and because of the wide range of forms that life may take. This chapter reviews observations of exoplanets and discusses astrobiological definitions of habitability and the likelihood of finding life beyond the Earth, both within and outside the Solar system.
@@ 1 Introduction
As of October 2023, these techniques have together revealed over 5500 exoplanets (NASA, 2023). The question if some of the planets within or beyond our Solar system harbour life goes back centuries, with recorded suggestions going back to e.g. Huygens (1698) and Herschel (1795). This paper reviews the status of the search for habitable worlds in the Universe from an astronomical point of view. Section 2 describes the observed properties of the known exoplanet population. Section 3 reviews the concept of planetary habitability, and estimates the number of potentially habitable worlds. Section 4 discusses limits to planetary habitability. Section 5 describes searches for life beyond the Earth. Section 6 draws conclusions and outlines future prospects in this field.
@@ 2 The observed exoplanet population
After 25 years of exoplanet research, it is clear that planets are more common than stars. On average, there are at least 1.1 planets per star, but the distribution depends on mass and composition. While only one in 10 stars has a Jupiter-size planet, 1 in 3 has a Neptune-sized planet, and every star (on average) has an Earth-sized planet. Stars with higher fractions of elements heavier than helium tend to have more gas giant planets, which is consistent with scenarios where planetary cores are built up from heavy elements, although the trend does not hold up for terrestrial planets (Buchhave et al., 2012).
@@
The searches for exoplanets have revealed several planet types that do not occur in the Solar system. In particular, most planets in the Universe turn out to be intermediate size, i.e. with a mass between that of Neptune and the Earth. Planets orbiting very close to their host stars are another example of planets without local analog. Gas giants cannot form close to their host stars, so the discovery of hot Jupiters indicates that planets may migrate to other orbits after their formation. This type of planet, while easily visible, is actually quite rare, with an occurrence rate below 1%. On the other hand, direct imaging has revealed planets very far from their host stars (several times as far as in the Solar system), and even planets without host stars, such as the Jupiter-sized free-floating planets in the Orion Nebula (Pearson and McCaughrean, 2023). Orphan planets may result from dynamical ejection from multiple systems, or from formation processes like those of low-mass stars.
The temperatures of exoplanets are essential input to know their physical state. Planetary temperatures can be roughly estimated by assuming an equilibrium between heating by starlight and cooling by thermal radiation. The temperature then mainly depends on the stellar luminosity and the orbital size, which are measured, and on planetary albedo and emissivity, for which fiducial values must be adopted. This method depends on the assumed rotation rate, and ignores the greenhouse effect (noticeable for Earth, strong on Venus) and heat from the planetary interior, both as leftover from the formation (important for Jupiter and Saturn), and from the decay of radioactive elements. The necessary input data to estimate these effects are usually not available for exoplanets, but reasonable assumptions can be made to construct models. Temperatures derived this way range from below freezing to above the melting points of rocks and metals. The warmest Earth-sized planets and the majority of detected intermediate-size planets are therefore thought to possess magma oceans on their surfaces, similar to the Earth in its early days. In the future, imaging of exoplanets at multiple infrared wavelengths will allow us to directly measure their emission temperatures.
@@
Studying planetary atmospheres is useful to learn about planetary (volatile) composition, as atmospheres are more accessible than surfaces or interiors, although atmospheric composition is not necessarily representative of planetary composition, as the example of the Earth shows. Studying atmospheres is also useful to learn about planetary formation (e.g., from dust or from gas) and subsequent evolution (e.g., migration to other orbits, or pollution by collisions with other bodies). Third, studying atmospheres gives clues to planetary habitability, since an atmosphere allows to retain volatile material (important for the circulation and transport of organics), by its condensation into clouds. In our Solar system, all planets and several moons possess atmospheres, but their properties vary widely. Jupiter and Saturn have thick atmospheres of Solar composition (mostly hydrogen and helium), indicating that they formed with the Solar system. The atmospheres of Uranus and Neptune, while also thick and hydrogen-helium dominated, are enriched in condensable gases (CO2, H2O, CH4), indicating that their formation ended when the Solar system lost its gas. In contrast, the rocky planets have thin heavy atmospheres, consisting of CO2 (Venus and Mars) or N2 (Earth; and also Saturn’s largest moon Titan). These planets lost their initial hydrogen-helium atmospheres (as evidenced by the depletion of heavy noble gases) and gained their current ’secondary’ atmospheres by outgassing from magma oceans (Kite and Barnett, 2020) and delivery by comets and asteroids (Broadley et al., 2022). After their formation, they evolved significantly: Mars and Venus lost their water, and the Earth atmosphere was enriched in O2 by biogenic processes. For a review of volatile segregation and processes during planet formation, see Suer et al. (2023).
@@
The atmospheres of exoplanets can be studied if the planet is seen in transit or in direct detection. For transiting planets, the atmospheric composition can be obtained by comparing the spectrum during transit (when the starlight is filtered by the planetary atmosphere) with that during eclipse, when the planet passes behind the star so that pure starlight is observed. Between transit and eclipse, a mix of direct and reflected starlight is observed in the visible part of the spectrum, and a mix of stellar and planetary emission in the infrared. By tracking this mix along the orbit (in a so-called phase curve), longitudinal structure can be observed, such as differences between the morning and evening sides of the planet. Such information can also be obtained from time-resolved transit observations, in particular during ingress and egress. For an overview of these techniques see Crossfield (2015) and Deming et al. (2019); for an overview of results see Madhusudhan (2019). About a dozen atomic and a dozen molecular species have been found prior to the launch of the James Webb Space Telescope (JWST) in December 2021. Highlights of JWST results in this area include the detection of SO2 in WASP-39b (a sign of active photochemistry; Tsai et al. 2023), and the simultaneous detection of CO2 and CH4 in K2-18b (Bell et al., 2023). So far such measurements are only possible for gas giants; for rocky planets, even JWST only obtains upper limits (Zieba et al., 2023; Greene et al., 2023). Quantitative analysis of exoplanet atmospheres remains difficult: estimated abundances depend strongly on the model for the background atmosphere.
@@ 3 Definitions of planetary habitability
Of the CHNOPS elements, only hydrogen has existed since the first three minutes of the Universe, when atoms formed out of subatomic particles. Elements heavier than lithium were formed later by nuclear fusion at the centers of stars; elements heavier than iron were formed in supernovae, neutron star mergers, and other energetic events leading to high neutron fluxes. In the so-called s- and r-processes, the capture of single resp. multiple neutrons is followed by $\beta$-decay into new elements. Building up sufficient CHNOPS to form rocky planets with water and biomolecules has taken $\sim$2 billion years (Dayal et al., 2015), which leaves $>$10 billion years for the development of life on planets.
@@
The third requirement for life as we know it is a solvent. Only in solution, chemistry proceeds fast enough to build up complex molecules such as proteins. The liquid acts as a transporter which brings organics together. For effective mixing, water flows are important, which call for temperature gradients and/or tidal forces from a satellite. The fourth and final basic requirement is the existence of stable benign conditions, which implies a stable orbit around a long-lived star without strong flares or other variability. The lifetimes of stars decrease rapidly with increasing mass, which renders stars significantly more massive than the Sun as inhospitable, as their lifetimes are less than a billion years. The most common type of stars in our Galaxy are red dwarfs with masses of $\sim$0.5 solar masses and lifetimes of $>$10 billion years. Many planets have been found around such stars, including our nearest neighbour Proxima b (Anglada-Escudé et al., 2016) and the well-known TRAPPIST-1 system with seven temperate Earth-sized planets (Gillon et al. 2017; Fig. 1 bottom). However, the high flaring activity of red dwarfs may be a limit to the development of life on the planets orbiting them.
@@
The number of (potentially) habitable exoplanets may be estimated from the properties of known exoplanets, by applying limits to their sizes (for atmospheric pressure) and equilibrium temperatures, so that liquid water may exist on their surfaces. An upper limit to the number is obtained by taking the freezing and boiling points of pure water (273 and 373 K) as temperature limits; we call these the ’naive’ limits. More realistic limits including climate and feedback effects would be temperatures between 246–295 K (the ’loose’ limits), or 260–289 K (the ’strict’ limits) (Kopparapu et al., 2013). As of August 2023, 5502 exoplanet detections have been confirmed (NASA, 2023), of which 378 have sizes between 0.4 and 2.5 Earth radii and temperatures within the ’naive’ limits, of which 147 fall within the ’loose’ limits, of which 98 within the ’strict’ limits. These numbers are approximate because the limits do not consider gases other than H2O, cloud effects, or 3D climate effects.
@@ 4 Bottlenecks for life on other planets
Even if our Galaxy contains millions of Earth-like planets, observations suggest that very few (if any) of these contain technologically advanced life forms. Searches for radio signals from extraterrestrial life have been conducted for $>$50 years without any positive result. In addition, there is no scientific evidence that alien civilizations have ever visited the Earth, which may be a stronger constraint (Grimaldi, 2017). The discrepancy between the number of habitable planets and the number of observed civilizations is known as the Fermi paradox, and suggests that evolution of life as it did on Earth is rare (although it is hard to say just how rare). This section discusses scenarios that aim to explain why this type of evolution may be rare.
One bottleneck for the emergence of life is that some planets may not have the right chemical conditions, even though they lie in the HZ of their host star (Krijt et al., 2022). Lack of surface water (or other suitable solvent) would be such a case, as the examples of Venus and Mars today indicate. Isotopic analysis of zircon minerals indicates that water has been present on the Earth’s surface for $\sim$4 billion years (Genda, 2016), which is longer than the oldest fossils, but shorter than the age of the Earth. Not much water is needed to start life: the Earth contains $\sim$30 oceans’ worth of water, of which 1 on the surface (Peslier et al., 2017; Hirose et al., 2021), which is only 0.023% by mass, much less than e.g. Europa or Enceladus. These bodies may actually have too much water: at depths greater than $\sim$100 km, the high pressure turns water into ice, which impedes chemical interaction with the ocean floor (Noack et al., 2016). Another case of chemical inhabitability would be a lack of atmospheric CO2 or CH4, which provide the greenhouse effect needed around faint young stars. Lack of nitrogen, phosphorus, or sulphur compounds may also inhibit the build-up of a fully functional biochemistry.
@@ 5 Searches for extraterrestrial life
This section discusses ongoing and planned searches for life on other planets. Several strategies exist for such searches, but most have in common that they seek to identify life by the relatively complex physical and/or chemical structures it produces, which are out of an equilibrium state. Some searches use the abundances of certain elements or isotopes, others use specific molecules or combinations of molecules. An example of the latter is the simultaneous presence of O2 and CH4, which would under abiotic circumstances react with each other. Their simultaneous presence in the Earth’s atmosphere is caused by their continuous production, respectively by cyanobacteria (as well as plants and trees) and by anaerobic methanogenic microbes (as well as ruminant animals). Searching for complex organic molecules is not enough: organics with $>$10 atoms have been found in interstellar gas clouds (McGuire, 2022), and amino acids have been found in both meteorites (such as Allende) and comets (e.g, 67P) which are clearly lifeless (Altwegg et al., 2019).
@@
For exoplanets, three types of biosignatures are often considered (Schwieterman et al., 2018). The first type are atmospheric gases such as O2, the main source of which on Earth is photosynthesis of H2O + CO2. The O2 is a byproduct of the synthesis of organic matter, and may be detected by its A-band, a spectroscopic feature in the far red part of the spectrum. This wavelength is accessible to ground-based telescopes, and the detection of O2 in exoplanet atmospheres is a science goal of the European Extremely Large Telescope which is starting operations in 2027 in Chile (Snellen et al., 2013). Other searches use O3, the photolytic byproduct of O2, which has a spectral signature at mid-infrared wavelengths, which may be detectable by the future Large Interferometer For Exoplanets (Quanz et al., 2022). The problem with using O2 as a biosignature is that it also has an abiotic formation route, namely the photodissociation of H2O by ultraviolet starlight, followed by the escape of the light H atoms from the atmosphere (Luger and Barnes, 2015; Schaefer et al., 2016). The heavier O atoms stay behind and react with OH into O2, which therefore cannot be considered an unambiguous biosignature. Furthermore, this signature would not have worked for the first 2.5 billion years (50%) of Earth history, when our atmosphere did not contain significant O2.
@@ 6 Conclusions and future prospects
The 2020s will see many new facilities, both in space and on the ground, which will make important contributions to this field. Launched in April 2023, the ESA mission JUICE will arrive in 2031 at the Jovian moon system for exploration of Europa, Ganymede, and Callisto. In 2024, NASA’s Clipper mission will follow, which is targeted toward Europa, and in 2028, Dragonfly will travel to Titan and explore its conditions with drone-type technology. Scheduled for launch in 2026, the PLATO mission will observe transits of Earth-like planets around Sun-like stars, and allow their basic characterization in terms of temperature and bulk density. The ARIEL mission will follow in 2028, which will observe transit spectra of mid-size exoplanets, allowing to assess the compositions of their atmospheres. Meanwhile, the ELT will have had first light in 2027, and started its searches for the O2 A-band in giant exoplanet atmospheres and rocky planets around red dwarfs. Direct imaging of exoplanets will become possible for smaller and older planets with the ELT from the ground and the NASA’s Roman telescope (to be launched 2027) from space. At the end of the decade, ESA’s Comet Interceptor (2029) will visit a pristine comet, which has never been close to the Sun before, and inform us about the physical and chemical conditions during the formation of the Solar system. In the next two decades, NASA’s Habitable Worlds Observatory (HWO) will take direct images of rocky exoplanets in the optical, and the Large Interferometer For Exoplanets (LIFE) may take their mid-infrared spectra. These latter missions will take significant steps in the characterization of rocky exoplanets, and the testing and development of proposed biosignatures.

### v2

@@
11institutetext: Floris van der Tak 22institutetext: Space Research Organization Netherlands (SRON) & Kapteyn Astronomical Institute, University of Groningen, Landleven 12, 9747 AD Groningen, The Netherlands; 22email: vdtak@sron.nl
Likelihood and appearance of life beyond the Earth: An astronomical perspective
Book chapter, to appear in First Contact: Aliens and Humans in Contemporary Science Fiction, eds. M. van Dijk, F. Bosman, K. Glimmerveen (Springer)
[…]
@@ 1 Introduction
As of October 2025, these techniques have together revealed over 6000 exoplanets (NASA, 2025). The question if some of the planets within or beyond our Solar system harbour life goes back centuries, with recorded suggestions going back to e.g. Huygens (1698) and Herschel (1795). This paper reviews the status of the search for habitable worlds in the Universe from an astronomical point of view. Section 2 describes the observed properties of the known exoplanet population. Section 3 reviews the concept of planetary habitability, and estimates the number of potentially habitable worlds. Section 4 discusses limits to planetary habitability. Section 5 describes searches for life beyond the Earth. Section 6 draws conclusions and outlines future prospects in this field.
@@ 2 The observed exoplanet population
After 30 years of exoplanet research, it is clear that planets are more common than stars. On average, there are at least 1.1 planets per star, but the distribution depends on mass and composition. While only one in 10 stars has a Jupiter-size planet, 1 in 3 has a Neptune-sized planet, and every star (on average) has an Earth-sized planet. Stars with higher fractions of elements heavier than helium tend to have more gas giant planets, which is consistent with scenarios where planetary cores are built up from heavy elements, although the trend does not hold up for terrestrial planets (Buchhave et al., 2012).
@@
The searches for exoplanets have revealed several planet types that do not occur in the Solar system. In particular, most planets in the Universe turn out to be intermediate size, i.e. with a mass between that of Neptune and the Earth. Planets orbiting very close to their host stars are another example of planets without local analog. Gas giants are unlikely to form close to their host stars, so the discovery of hot Jupiters indicates that planets may migrate to other orbits after their formation. This type of planet, while easily visible, is actually quite rare, with an occurrence rate below 1%. On the other hand, direct imaging has revealed planets very far from their host stars (several times as far as in the Solar system), and even planets without host stars, such as the Jupiter-sized free-floating planets in Upper Sco (Miret-Roig et al., 2022). Orphan planets may result from dynamical ejection from multiple systems, or from formation processes like those of low-mass stars.
The temperatures of exoplanets are essential input to know their physical state. Planetary temperatures can be roughly estimated by assuming an equilibrium between heating by starlight and cooling by thermal radiation. The temperature then mainly depends on the stellar luminosity and the orbital size, which are measured, and on planetary albedo and emissivity, for which fiducial values must be adopted. This method depends on the assumed rotation rate, and ignores the greenhouse effect (noticeable for Earth, strong on Venus) and heat from the planetary interior, both as leftover from the formation (important for Jupiter and Saturn), and from the decay of radioactive elements. The necessary input data to estimate these effects are usually not available for exoplanets, but reasonable assumptions can be made to construct models. Temperatures derived this way range from below the boiling point of nitrogen to above the melting points of rocks and metals. The warmest Earth-sized planets and the majority of detected intermediate-size planets are therefore thought to possess magma oceans on their surfaces, similar to the Earth in its early days. In the future, imaging of exoplanets at multiple infrared wavelengths will allow us to directly measure their emission temperatures.
@@
Studying planetary atmospheres is useful to learn about planetary (volatile) composition, as atmospheres are more accessible than surfaces or interiors, although atmospheric composition is not necessarily representative of planetary composition, as the example of the Earth shows. Studying atmospheres is also useful to learn about planetary formation (e.g., from dust or from gas) and subsequent evolution (e.g., migration to other orbits, or pollution by collisions with other bodies). Third, studying atmospheres gives clues to planetary habitability, since an atmosphere allows to retain volatile material (important for the circulation and transport of organics), by its condensation into clouds. In our Solar system, all planets and several moons possess atmospheres, but their properties vary widely. Jupiter and Saturn have thick atmospheres of Solar composition (mostly hydrogen and helium), indicating that they formed with the Solar system. The atmospheres of Uranus and Neptune, while also thick and hydrogen-helium containing, consist mostly of condensable gases (NH3, H2O, CH4), indicating that their formation ended when the Solar system lost its gas. In contrast, the rocky planets have thin heavy atmospheres, consisting of CO2 (Venus and Mars) or N2 (Earth; and also Saturn’s largest moon Titan). These planets lost their initial hydrogen-helium atmospheres (as evidenced by the depletion of heavy noble gases) and gained their current ’secondary’ atmospheres by outgassing from magma oceans (Kite and Barnett, 2020) and delivery by comets and asteroids (Broadley et al., 2022). After their formation, they evolved significantly: Mars and Venus lost their water, and the Earth atmosphere was enriched in O2 by biogenic processes. For a review of volatile segregation and processes during planet formation, see Suer et al. (2023).
@@
The atmospheres of exoplanets can be studied if the planet is seen in transit or in direct detection. For transiting planets, the atmospheric composition can be obtained by comparing the spectrum during transit (when the starlight is filtered by the planetary atmosphere) with that during eclipse, when the planet passes behind the star so that pure starlight is observed. Between transit and eclipse, a mix of direct and reflected starlight is observed in the visible part of the spectrum, and a mix of stellar and planetary emission in the infrared. By tracking this mix along the orbit (in a so-called phase curve), longitudinal structure can be observed, such as differences between the morning and evening sides of the planet. Such information can also be obtained from time-resolved transit observations, in particular during ingress and egress. For an overview of these techniques see Crossfield (2015) and Deming et al. (2019); for an overview of results see Madhusudhan (2019). About a dozen atomic and a dozen molecular species have been found prior to the launch of the James Webb Space Telescope (JWST) in December 2021. Highlights of JWST results in this area include the detection of SO2 in WASP-39b (a sign of active photochemistry; Tsai et al. 2023), and the simultaneous detection of CO2 and CH4 in K2-18b (Bell et al., 2023). So far such measurements are only possible for gas giants, although for lower-mass planets, first results are appearing (Hu et al., 2024; Ducrot et al., 2025). Quantitative analysis of exoplanet atmospheres remains difficult: estimated abundances depend strongly on the model for the background atmosphere.
@@ 3 Definitions of planetary habitability
Of the CHNOPS elements, only hydrogen has existed since the first three minutes of the Universe, when atoms formed out of subatomic particles. Elements heavier than lithium were formed later by nuclear fusion at the centers of stars; elements heavier than iron were formed in supernovae, neutron star mergers, and other energetic events leading to high neutron fluxes. In the so-called s- and r-processes, the capture of single resp. multiple neutrons is followed by $\beta$-decay into new elements. Building up sufficient CHNOPS to form rocky planets with water and biomolecules has taken $\sim$2 billion years (Dayal et al., 2015), which leaves >>10 billion years for the development of life on planets.
@@
The third requirement for life as we know it is a solvent. Only in solution, chemistry proceeds fast enough to build up complex molecules such as proteins. The liquid acts as a transporter which brings organics together. For effective mixing, water flows are important, which call for temperature gradients and/or tidal forces from a satellite. The fourth and final basic requirement is the existence of stable benign conditions, which implies a stable orbit around a long-lived star without strong flares or other variability. The lifetimes of stars decrease rapidly with increasing mass, which renders stars significantly more massive than the Sun as inhospitable, as their lifetimes are less than a billion years. The most common type of stars in our Galaxy are red dwarfs with masses of $\sim$0.5 solar masses and lifetimes of >>10 billion years. Many planets have been found around such stars, including our nearest neighbour Proxima b (Anglada-Escudé et al., 2016) and the well-known TRAPPIST-1 system with seven temperate Earth-sized planets (Gillon et al. 2017; Fig. 1 bottom). However, the high flaring activity of red dwarfs may be a limit to the development of life on the planets orbiting them.
@@
The number of (potentially) habitable exoplanets may be estimated from the properties of known exoplanets, by applying limits to their sizes (for atmospheric pressure) and equilibrium temperatures, so that liquid water may exist on their surfaces. An upper limit to the number is obtained by taking the freezing and boiling points of pure water (273 and 373 K) as temperature limits; we call these the ’naive’ limits. More realistic limits including climate and feedback effects would be temperatures between 246–295 K (the ’loose’ limits), or 260–289 K (the ’strict’ limits) (Kopparapu et al., 2013). As of August 2023, 5502 exoplanet detections have been confirmed (NASA, 2025), of which 378 have sizes between 0.4 and 2.5 Earth radii and temperatures within the ’naive’ limits, of which 147 fall within the ’loose’ limits, of which 98 within the ’strict’ limits. These numbers are approximate because the limits do not consider gases other than H2O, cloud effects, or 3D climate effects.
@@ 4 Bottlenecks for life on other planets
Even if our Galaxy contains millions of Earth-like planets, observations suggest that very few (if any) of these contain technologically advanced life forms. Searches for radio signals from extraterrestrial life have been conducted for >>50 years without any positive result. In addition, there is no scientific evidence that alien civilizations have ever visited the Earth, which may be a stronger constraint (Grimaldi, 2017). The discrepancy between the number of habitable planets and the number of observed civilizations is known as the Fermi paradox, and suggests that evolution of life as it did on Earth is rare (although it is hard to say just how rare). This section discusses scenarios that aim to explain why this type of evolution may be rare.
One bottleneck for the emergence of life is that some planets may not have the right chemical conditions, even though they lie in the HZ of their host star (Krijt et al., 2022). Lack of surface water (or other suitable solvent) would be such a case, as the examples of Venus and Mars today indicate. Isotopic analysis of zircon minerals indicates that water has been present on the Earth’s surface for $\sim$4 billion years (Genda, 2016), which is longer than the oldest fossils, but shorter than the age of the Earth. Not much water is needed to start life: the Earth contains $\sim$30 oceans’ worth of water, of which 1 on the surface (Peslier et al., 2017; Hirose et al., 2021), which is only 0.023% by mass, much less than e.g. Europa or Enceladus. These bodies may actually have too much water: at ocean depths greater than $\sim$100 km, the high pressure turns water into ice, which impedes chemical interaction with the ocean floor (Noack et al., 2016). Another case of chemical inhabitability would be a lack of atmospheric CO2 or CH4, which provide the greenhouse effect needed around faint young stars. Lack of nitrogen, phosphorus, or sulphur compounds may also inhibit the build-up of a fully functional biochemistry.
@@ 5 Searches for extraterrestrial life
This section discusses ongoing and planned searches for life on other planets. Several strategies exist for such searches, but most have in common that they seek to identify life by the relatively complex physical and/or chemical structures it produces, which are out of an equilibrium state. Some searches use the abundances of certain elements or isotopes, others use specific molecules or combinations of molecules. An example of the latter is the simultaneous presence of O2 and CH4, which would under abiotic circumstances react with each other. Their simultaneous presence in the Earth’s atmosphere is caused by their continuous production, respectively by cyanobacteria (as well as plants and trees) and by anaerobic methanogenic microbes (as well as ruminant animals). Searching for complex organic molecules is not enough: organics with >>10 atoms have been found in interstellar gas clouds (McGuire, 2022), and amino acids have been found in both meteorites (such as Allende) and comets (e.g, 67P) which are clearly lifeless (Altwegg et al., 2019).
@@
For exoplanets, three types of biosignatures are often considered (Schwieterman et al., 2018). The first type are atmospheric gases such as O2, the main source of which on Earth is photosynthesis of H2O + CO2. The O2 is a byproduct of the synthesis of organic matter, and may be detected by its A-band, a spectroscopic feature in the far red part of the spectrum. This wavelength is accessible to ground-based telescopes, and the detection of O2 in exoplanet atmospheres is a science goal of the European Extremely Large Telescope which is starting operations in 2029 in Chile (Snellen et al., 2013). Other searches use O3, the photolytic byproduct of O2, which has a spectral signature at mid-infrared wavelengths, which may be detectable by the future Large Interferometer For Exoplanets (Quanz et al., 2022). The problem with using O2 as a biosignature is that it also has an abiotic formation route, namely the photodissociation of H2O by ultraviolet starlight, followed by the escape of the light H atoms from the atmosphere (Luger and Barnes, 2015; Schaefer et al., 2016). The heavier O atoms stay behind and react with OH into O2, which therefore cannot be considered an unambiguous biosignature. Furthermore, this signature would not have worked for the first 2.5 billion years (50%) of Earth history, when our atmosphere did not contain significant O2.
@@ 6 Conclusions and future prospects
The 2020s will see many new facilities, both in space and on the ground, which will make important contributions to this field. Launched in April 2023, the ESA mission JUICE will arrive in 2031 at the Jovian moon system for exploration of Europa, Ganymede, and Callisto. In 2024, NASA’s Clipper mission has followed, which is targeted toward Europa, and in 2028, Dragonfly will travel to Titan and explore its conditions with drone-type technology. Scheduled for launch in 2026, the PLATO mission will observe transits of Earth-like planets around Sun-like stars, and allow their basic characterization in terms of temperature and bulk density. The ARIEL mission will follow in 2028, which will observe transit spectra of mid-size exoplanets, allowing to assess the compositions of their atmospheres. Meanwhile, the ELT will see its first light in 2029, and started its searches for the O2 A-band in giant exoplanet atmospheres and rocky planets around red dwarfs. Direct imaging of exoplanets will become possible for smaller and older planets with the ELT from the ground and the NASA’s Roman telescope (to be launched 2027) from space. At the end of the decade, ESA’s Comet Interceptor (2029) will visit a pristine comet, which has never been close to the Sun before, and inform us about the physical and chemical conditions during the formation of the Solar system. In the next two decades, NASA’s Habitable Worlds Observatory (HWO) will take direct images of rocky exoplanets in the optical, and the Large Interferometer For Exoplanets (LIFE) may take their mid-infrared spectra. These latter missions will take significant steps in the characterization of rocky exoplanets, and the testing and development of proposed biosignatures.